OpenStudy (anonymous):

PLEASE HELP ME DIVIDE THIS RATIONAL EXPRESSION 4C-12 ------ 4C+8 ------ C-3 ----- C^2-4

OpenStudy (anonymous):

Well you have ((4c-12)/(4c+8))/((c-3)/(c^2-4)) So reciprocate. Giving: ((4c-12)(c^2-4))/((4c+8)(c-3)) Distribute. Then divide.

OpenStudy (anonymous):

Is that: \[\frac{\frac{4c - 12}{4c+8}}{\frac{c - 3}{c^{2}-4}}\] ??? cant tell for sure which way the fractions are set up. could be that its: \[\frac{\frac{\frac{4c-12}{4c+8}}{c-3}}{c^{2}-4}\] or some such

OpenStudy (anonymous):

YES ITS THE FIRST ONE

OpenStudy (anonymous):

CAPS LOCK FTW?!?!

OpenStudy (anonymous):

Lol, well in that case just do what malevolence posted first. Dividing by a fraction is the same as multiplying by the fractions reciprocal, remember?

OpenStudy (anonymous):

yeah learned this a long time ago, and really dont remember any of it